OpenStudy (anonymous):

1. A car drives 12 miles north, turns, and then drives 8 miles east. How far is the car from its starting point? a. Draw a right triangle to represent this situation. Label the lengths of the known sides. Label the length of the unknown side x. b. Use the Pythagorean theorem to find the length of the missing side. Be sure to indicate which root represents the answer and write the answer in simplest radical form. c. Use a calculator to change the answer to decimal form. Round the answer to the nearest tenth.

OpenStudy (anonymous):

Using this decimal number, write a sentence to explain the meaning of the answer in this situation

OpenStudy (anonymous):

@dg2 @ganeshie8

OpenStudy (aum):



OpenStudy (anonymous):

how do i do b & c?

OpenStudy (aum):

b) Use Pythagorean theorem: AB^2 + BC^2 = AC^2 12^2 + 8^2 = x^2 144 + 64 = x^2 208 = x^2 \(x = \sqrt{208} = \sqrt{16*13} = 4\sqrt{13} \)

OpenStudy (anonymous):

thanks:)

OpenStudy (aum):

For b) write x = \(\pm 4\sqrt{13}\) and then choose the positive root: x = \(+4\sqrt{13}\)

OpenStudy (anonymous):

okay.

OpenStudy (aum):

c) Use your calculator to find \(+4\sqrt{13} = 14.4\)

OpenStudy (anonymous):

alright i will